OpenStudy (anonymous):

Nina was curious about the height of the Eiffel Tower. She used a 1.2 m model of the tower and measured its shadow at 2 p.m. The length of the shadow was 0.9 m. She then measured the Eiffel Tower’s shadow and found that it was 240 m. What is the height of the Eiffel Tower?

OpenStudy (anonymous):

you can solve \[\frac{1.2}{.9}=\frac{\text{height}}{240}\] via \[\text{height}=\frac{240\times 1.2}{.9}\]